Suppose that the thickness of a part used in a semiconductor is its critical dimension and that measurements of the thickness of a random sample of 18 such parts have the variance S2 = 0.68, where the measurements are in thousandths of an inch. The process is considered to be under control if the variation of the thickness is given by a variance not greater than 0.36. Assuming that the measurements constitute a random sample from a normal population, test the claim at the α =0.05 significance level.




1
Expert's answer
2021-12-14T08:12:29-0500

The hypotheses to be tested are,

"H_0: \sigma^2 =0.36\space vs\space H_A: \sigma^2 > 0.36"

"n=18,\space s^2=0.68"

The test statistic is,

"\chi^2 =\frac{(n-1) \times s^2}{\sigma^2}"


"\chi^2 = \frac{17 \times 0.68}{0.36} = 32.111"

The table value at "\alpha=0.05" with "n-1=18-1=17" degrees of freedom is, "\chi^2_{0.05,17}=27.8571" and the null hypothesis is rejected if "\chi^2\gt\chi^2_{0.05,17}".

Since "\chi^2=32.111\gt\chi^2_{0.05,17}=27.8571," we reject the null hypothesis and we conclude that there is sufficient evidence to show that the population variance is greater than 0.36 level at 5% level of significance and the process is out of control .